OpenStudy (anonymous):

What is a polynomial function in standard form with zeroes 1,2,3 and -1

OpenStudy (solomonzelman):

Okay a polynomial function f(x) with zeros \(\large\color{black}{ a }\), \(\large\color{black}{ b }\), \(\large\color{black}{ c }\), and \(\large\color{black}{ d}\), can be written as \(\large\color{black}{ f(x)=(x-a)(x-b)(x-c)(x-d) }\) The reason for this is a zero product property, because when one of the parenthesis is equal to zero, then the entire f(x) is also equal to zero.

OpenStudy (campbell_st):

well use the information is x = 1 is a zero then factor is x - 1 repeat the process for all the zeros x = 2 is a zero then x - 2 is a factor so P(x) = (x -1)(x -2)? ?? you need to find the last 2 binomial factors then expand it all hope it helps